# On the semi-additivity of the $1/2$-symmetric caloric capacity in the plane

### v3

@@ 2. Notation and basic definitions
$\gamma_{\text{op}}(E)=\gamma_{\Theta^{1/2},\text{op}}(E)$ $:=\sup\big{\{}\mu(E)\;:\;\mu\in\Sigma(E),\;\|\pazocal{P}_{\mu}\|_{L^{2}(\mu)\to L^{2}(\mu)}\leq 1\big{\}},$
$\gamma_{\text{sy},\text{op}}(E)$ $:=\sup\big{\{}\mu(E)\;:\;\mu\in\Sigma(E),\;\|\pazocal{P}_{\text{sy},\mu}\|_{L^{2}(\mu)\to L^{2}(\mu)}\leq 1\big{\}},$
@@
From the results of [MPr] and [He] one can deduce an important property of the above different capacities: they are all comparable.
@@
$\approx\gamma^{\prime}_{+}(E)\approx\overline{\gamma}^{\prime}_{+}(E)\approx\widetilde{\gamma}^{\prime}_{+}(E)\approx\gamma_{\text{{sy}},+}^{\prime}(E)$
$\hskip 134.5814pt\approx\gamma_{\text{{op}}}(E)\approx\gamma_{\text{{sy,op}}}(E).$
@@
To be precise, the capacities $\overline{\gamma}^{\prime}_{+},\gamma_{\text{{sy}},+}^{\prime}$ and $\gamma_{\text{{sy,op}}}$ are not introduced in the aforementioned references. To justify the comparability of the capacity $\gamma_{\text{{sy}},+}^{\prime}$ to all the rest, just observe that by the nonnegativity of $P_{\text{{sy}}}$ it is clear that
$\widetilde{\gamma}^{\prime}_{+}(E)\leq\gamma_{\text{sy},+}^{\prime}(E)\leq 2\widetilde{\gamma}^{\prime}_{+}(E).$
To do the same with $\overline{\gamma}^{\prime}_{+}$, we use that $P^{\ast}$ is the fundamental solution of the conjugate operator $\overline{\Theta}^{1/2}$. By an analogous proof to [MPr, Lemma 3.4] one deduces that admissible measures for $\overline{\gamma}_{+}$ present $n$-growth with respect to a dimensional constant. Therefore [He, Lemma 3.10 & Theorem 3.11] admit an analogous version for $P^{\ast}$, implying
$\overline{\gamma}_{+}(E)\lesssim\overline{\gamma}^{\prime}_{+}(E)\lesssim\gamma_{\text{{op}}}(E)\approx\widetilde{\gamma}_{+}\leq\overline{\gamma}_{+}(E).$
Hence, $\overline{\gamma}^{\prime}_{+}$ is also comparable to all the capacities of Theorem 2.1. Finally, to deduce the comparability of $\gamma_{\text{{sy,op}}}$ it suffices to check $\gamma_{\text{sy,op}}\lesssim\gamma_{\text{op}}$, since the reverse inequality can be verified to hold easily. Fix $\mu$ any admissible measure for $\gamma_{\text{{sy,op}}}(E)$ with $\gamma_{\text{{sy,op}}}(E)\leq 2\mu(E)$. Notice that since $\pazocal{P}_{\mu}$ and $\pazocal{P}^{\ast}_{\mu}$ are $n$-dimensional C-Z operators, then $\pazocal{P}_{\text{sy},\mu}$ also satisfies such property. Therefore, following an analogous proof to [To3, Theorem 2.16] one deduces, from the $L^{2}(\mu)$-boundedness of $\pazocal{P}_{\text{sy},\mu}$, its boundedness as an operator from the space of Borel finite signed measures $M(\mathbb{R}^{n+1})$ to $L^{1,\infty}(\mu)$. That is, there exists a constant $C>0$ such that for any $\nu\in M(\mathbb{R}^{n+1})$,
@@ 3. Even more capacities
for all the pair of points where such expression makes sense. Notice that it is also a nonnegative kernel and differentiable, with respect to $\overline{x}$ or $\overline{y}$, out of a set of null Lebesgue measure. We may also define $P_{\Lambda}^{\ast}$ and $P_{\text{sy},\Lambda}$ in an analogous way. In fact, for the sake of notation, we shall prove three basic properties only for $P_{\Lambda}$, but they can be checked to hold also for $P_{\Lambda}^{\ast}$ and $P_{\text{sy},\Lambda}$ with almost exactly the same proofs. Essentially the convolution kernels $P_{\Lambda},P_{\Lambda}^{\ast}$ and $P_{\text{sy},\Lambda}$ in $\mathbb{R}^{n+1}$ define $n$-dimensional C-Z operators. In fact, such suppressed kernels satisfy 3 as an additional property.
@@
Let $\Lambda:\mathbb{R}^{n+1}\to\mathbb{[}0,\infty)$ be a 1-Lipschitz function. Then, kernels $P_{\Lambda},P_{\Lambda}^{\ast}$ and $P_{\text{{sy}},\Lambda}$ define $n$-dimensional C-Z operators which satisfy being “well suppressed” at the points where $\Lambda>0$. Namely, if we fix any $\overline{x}\neq\overline{y}$ points in $\mathbb{R}^{n+1}$, the precise estimates that $P_{\Lambda}$ satisfies ((((as well as $P_{\Lambda}^{\ast}$ and $P_{\text{{sy}},\Lambda})$ are
@@
$\gamma_{\text{sy},2}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma(E),\,\|P_{\text{{sy}}}\ast\mu\|_{L^{2}(\mu)}\leq\mu(E)^{1/2}\big{\}}.$
It is clear that $\gamma_{\text{sy},+}^{\prime}(E)\leq\gamma_{\text{sy},2}(E)$. The fundamental lemma that will imply the reverse inequality is the following:
@@
Due to the above Lemma 3.2, we may add $\gamma_{\text{sy},2}$ to the statement of Theorem 2.1. In fact, defining
$\widetilde{\gamma}_{2}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma(E),\,\|P\ast\mu\|_{L^{2}(\mu)}\leq\mu(E)^{1/2},\|P^{\ast}\ast\mu\|_{L^{2}(\mu)}\leq\mu(E)^{1/2}\big{\}},$
it is clear that $\widetilde{\gamma}_{2}(E)\leq\gamma_{\text{sy},2}(E)$ and that $\widetilde{\gamma}^{\prime}_{+}(E)\leq\widetilde{\gamma}_{2}(E)$. Meaning that $\widetilde{\gamma}_{2}$ can be also added to Theorem 2.1.
@@ 3.2. Normalization by a uniform bound at all points of the support
$\gamma^{\star}_{+}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma(E),\,|P\ast\mu(\overline{x})|\leq 1,\,\forall\overline{x}\in E\big{\}},$
@@
Indeed, given $\mu\in\Sigma(E)$ admissible for $\gamma_{\text{op}}(E)$ with $\gamma_{\text{op}}(E)\leq 2\mu(E)$, proceeding as in Remark 2.2 we can find a function $h:E\to[0,1]$ such that $\int_{E}h\text{d}\mu\geq\mu(E)/2$ and $\|\pazocal{P}(h\mu)\|_{\infty}\leq C$, for some $C>0$ constant. We know that the latter estimate implies $\|\pazocal{P}(h\mu)\|_{L^{\infty}(\mu)}\leq C^{\prime}$ for some other constant $C^{\prime}>0$ (by a Cotlar type inequality analogous to that of [MaP, Lemma 5.4]). Applying now Cotlar’s inequality of [To3, Theorem 2.18], for example, we have
@@ 3.3. An equivalent variational capacity
We shall consider a construction already presented in the proof of Lemma 3.2: let $\psi$ be a radial test function with $0\leq\psi\leq 1$, $\psi\equiv 0$ in $B_{1/2}(0)$, $\psi\equiv 1$ in $\mathbb{R}^{n+1}\setminus{B_{1}(0)}$ and such that $\|\nabla\psi\|_{\infty}\leq 1$. Fix $E\subset\mathbb{R}^{n+1}$ compact set and $\mu\in\Sigma(E)$. For each $\tau>0$, write $\psi_{\tau}(\cdot):=\psi(\cdot/\tau)$ and define, for each $\overline{x}\in\mathbb{R}^{n+1}$ and $f\in L^{1}_{\text{loc}}(\mu)$.
@@
Let us prove, using [MPr, Lemma 2.1] (also valid for $P^{\ast}$ and $P_{\text{{sy}}}$) , that the regularized (continuous) kernel $P_{\text{{sy}}}\psi_{\tau}$ defines a C-Z convolution operator with constants not depending on $\tau$. It is clear that $|P_{\text{{sy}}}\psi_{\tau}(\overline{x})|\leq|\overline{x}|^{-n}$. Then, by the symmetry of $P_{\text{{sy}}}\psi_{\tau}$, it suffices to check that for any $\overline{x},\overline{x}^{\prime}\in\mathbb{R}^{n+1}$ with $\overline{x}\neq 0$ and $|\overline{x}-\overline{x}^{\prime}|\leq|\overline{x}|/2$,
@@
$|\displaystyle\big{\rvert}P_{\text{{sy}}}\psi_{\tau}(\overline{x})-P_{\text{{sy}}}\psi_{\tau}(\overline{x}^{\prime})\big{\rvert}$$\leq\big{\rvert}P_{\text{{sy}}}(\overline{x})-P_{\text{{sy}}}(\overline{x}^{\prime})\big{\rvert}\psi_{\tau}(\overline{x}^{\prime})+\big{\rvert}\psi_{\tau}(\overline{x})-\psi_{\tau}(\overline{x}^{\prime})\big{\rvert}P_{\text{{sy}}}(\overline{x})$
@@
$|\displaystyle\big{\rvert}P_{\text{{sy}}}\psi_{\tau}(\overline{x})-P_{\text{{sy}}}\psi_{\tau}(\overline{x}^{\prime})\big{\rvert}$$\leq\big{\rvert}P_{\text{{sy}}}(\overline{x})-P_{\text{{sy}}}(\overline{x}^{\prime})\big{\rvert}\psi_{\tau}(\overline{x}^{\prime})+\big{\rvert}1-\psi_{\tau}(\overline{x}^{\prime})\big{\rvert}P_{\text{{sy}}}(\overline{x})$
@@
$\gamma^{\star}_{\text{sy},\psi_{\tau},+}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma(E),\,|P_{\text{{sy}}}\psi_{\tau}\ast\mu(\overline{x})|\leq 1,\,\forall\overline{x}\in E\big{\}}.$
@@
We begin by proving 1. Let $(\tau_{k})_{k}$ be a monotonically decreasing sequence to 0, and let $\mu_{k}\in\Sigma(E)$ be admissible for $\gamma^{\star}_{\text{{sy}},\psi_{\tau_{k}},+}(E)$ and such that $\gamma^{\star}_{\text{{sy}},\psi_{\tau_{k}},+}(E)\leq 2\mu_{k}(E)$. Passing to a subsequence if necessary, assume $\mu_{k}\rightharpoonup\mu_{0}$ weakly [Ma, Theorem 1.23], i.e.
@@
It is not hard to prove that $\mu_{0}\in\Sigma(E)$. Now fix $m>0$ integer and $\tau\in[\tau_{m+1},\tau_{m}]$, and also assume $k>m$. By (3.7) we get that for $\overline{x}\in E$,
@@
and 1 follows. To prove 2, take $\mu$ admissible for $\gamma^{\star}_{\text{sy},+}(E)$. Using (3.7) we deduce
@@
implying that $\gamma^{\star}_{\text{sy},\psi_{\tau},+}(E)\geq\frac{\mu(E)}{1+C}$, and by the arbitrariness of $\mu$ we have
$\inf_{\tau>0}\gamma^{\star}_{\text{sy},\psi_{\tau},+}(E)\gtrsim\gamma^{\star}_{\text{sy},+}(E),$
which implies the desired result. ∎
@@
$\gamma_{\text{sy},\psi_{\tau},+}^{\star}(E)\underset{(1)}{\leq}\gamma_{\text{sy},\psi_{\tau},+}^{\prime}(E)\underset{(2)}{\leq}\gamma_{\text{sy},\psi_{\tau},2}(E)\underset{(3)}{\lesssim}\gamma_{\text{sy},\psi_{\tau},\text{op}}(E)\underset{(4)}{\lesssim}\gamma_{\text{sy},\psi_{\tau},+}(E)\underset{(5)}{\leq}\gamma_{\text{sy},\psi_{\tau},+}^{\star}(E).$
Inequalities (1) and (2) are trivial. To verify (3), proceed, for example, as in the proof of $\gamma_{\text{sy},2}(E)\lesssim\gamma_{\text{sy},\text{op}}(E)$ in Lemma 3.2 (although the arguments could be possibly simplified using the continuity of $P_{\text{{sy}}}\psi_{\tau}$). The proof of (4) goes as in Remark 2.2, and it relies on [To3, Theorem 2.16] and [Ch, Ch.VII, Theorem 23]. it is important to notice that a fundamental property that ensures the validity of (3) and (4) is that the C-Z constants of $P_{\text{{sy}}}\psi_{\tau}$ are independent of $\tau$. Finally, (5) holds by the continuity of $\pazocal{P}_{\text{sy},\psi_{\tau}}\mu$ for each $\tau>0$, that has already been argued in the proof of Lemma 3.2. Observe that the previous chain of estimates also holds changing every $\gamma_{\text{sy}}$ for $\widetilde{\gamma}$, where we ask the respective normalization conditions in each case for both kernels $P\psi_{\tau}$ and $P^{\ast}\psi_{\tau}$.
@@
Now, bearing in mind that we have found
  • $\star$
@@
for some positive constants $C_{1},C_{2}$, we define the variational capacity:
Definition 3.1 (Operator $\pazocal{S}$ and variational capacity).
Let $E\subset\mathbb{R}^{n+1}$ be a compact set and choose $\tau_{0}$ small enough so that
$\gamma_{\text{{sy}},\psi_{\tau_{0}},2}(E)\leq 2C_{1}C_{2}\,\gamma_{\text{sy},\text{op}}(E).$
Let $\pazocal{S}$ be the convolution operator associated to the kernel $P_{\text{{sy}}}\psi_{\tau_{0}}$ (that depends on $\tau_{0}$ and thus on $E$). We define the variational capacity of $E$ as
$\gamma_{\text{var}}(E):=\sup\Bigg{\{}\frac{\mu(E)^{2}}{\mu(E)+\int_{E}|\pazocal{S}\mu|^{2}\mathop{}\!\mathrm{d}\mu}\,:\,\mu\in\Sigma(E)\Bigg{\}}.$
@@
For each $\mu\in\Sigma(E)$ define
@@
As it has already been mentioned in the proof of Lemma 3.3, it is not difficult to prove that $\Sigma(E)$ is closed and sequentially compact with respect to weak convergence of measures. In fact, $\Sigma(E)$ is compact (since it is contained in the space of finite signed Radon measures on $E$, which is metrizable, thought as the dual of the separable space $\pazocal{C}_{c}(\mathbb{R}^{n+1})$). Moreover, it is clear that if $\mu_{k}\rightharpoonup\mu_{0}$ on $\Sigma(E)$, then $\mu_{k}(E)\to\mu_{0}(E)$. In this setting, we also claim that
@@
[…]
@@
Further, observe that we can restrict ourselves to the case that there exists $\mu\in\Sigma(E)$ such that $F(\mu)>0$. If this was not the case, since $\int_{E}|\pazocal{S}\mu|\mathop{}\!\mathrm{d}\mu\lesssim\text{diam}(E)^{2n}/\tau_{0}^{n}<\infty$, we would have, necessarily, $\mu(E)=0$, implying, by definition, $\gamma_{\text{var}}(E)=0=\gamma_{\text{{sy}},\psi_{\tau_{0}},2}(E)$.
So let us assume that there exists $\mu\in\Sigma(E)$ such that $F(\mu)>0$, meaning $\gamma_{\text{var}}(E)>0$ Let us begin by proving that any extremal $\mu_{0}$ for $F$ in this setting satisfies
@@
If this was not the case, then $\int_{E}|\pazocal{S}\mu_{0}|^{2}\mathop{}\!\mathrm{d}\mu_{0}=M\mu_{0}(E)$ for some $M>1$. Define $\mu_{1}:=M^{-1/2}\mu_{0}\in\Sigma(E)$ and notice that
$F(\mu_{1})$ $=\frac{M^{-1}\mu_{0}(E)^{2}}{M^{-1/2}\mu_{0}(E)+\int_{E}|\pazocal{S}\mu_{1}|^{2}\mathop{}\!\mathrm{d}\mu_{1}}$
$=\frac{M^{-1}\mu_{0}(E)^{2}}{M^{-1/2}\mu_{0}(E)+M^{-1/2}\mu_{0}(E)}=\frac{\mu_{0}(E)}{2M^{1/2}}>\frac{\mu_{0}(E)}{1+M}=F(\mu_{0}),$
@@
Let us notice that (3.10) holds in general for any extremal measure $\mu_{0}$ for $F$. Indeed, in the case that for any $\mu\in\Sigma(E)$ one had $F(\mu)=0$, then $\mu(E)=0$. So any measure would be extremal and (3.10) would hold trivially.
Finally we give a new version of Theorem 2.1 that relates the capacities presented in this section:
@@
All of the above capacities defined via positive Borel measures are comparable. More precisely, for $E\subset\mathbb{R}^{n+1}$ compact,
$\gamma_{+}(E)\approx\gamma^{\prime}_{+}(E)\approx\gamma^{\star}_{+}(E),$
where the respective bounds of the potentials can be taken indifferently and independently for $P,P^{\ast},P_{\text{{sy}}}$ or both $P$ and $P^{\ast}$ simultaneously. Moreover, they are also comparable to
$\widetilde{\gamma}_{2}(E),\;\gamma_{\text{{sy}},2}(E),\;\gamma_{\text{{op}}}(E),\;\gamma_{\text{{sy,op}}}(E)\quad\text{and}\quad\gamma_{\text{{var}}}(E).$

### v4

@@ 2. Notation and basic definitions
$\gamma_{\text{op}}(E)=\gamma_{\Theta^{1/2},\text{op}}(E)$ $:=\sup\big{\{}\mu(E)\;:\;\mu\in\Sigma_{n}(E),\;\|\pazocal{P}_{\mu}\|_{L^{2}(\mu)\to L^{2}(\mu)}\leq 1\big{\}},$
$\gamma_{\text{sy},\text{op}}(E)$ $:=\sup\big{\{}\mu(E)\;:\;\mu\in\Sigma_{n}(E),\;\|\pazocal{P}_{\text{sy},\mu}\|_{L^{2}(\mu)\to L^{2}(\mu)}\leq 1\big{\}},$
@@
From the results of [MPr] and [He] one can deduce an important property of the above different capacities: most of them are comparable.
@@
$\approx\gamma_{\text{{op}}}(E)\approx\gamma_{\text{{sy,op}}}(E).$
@@
To be precise, the capacity $\gamma_{\text{{sy,op}}}$ is not introduced in the aforementioned references. To deduce the comparability of $\gamma_{\text{{sy,op}}}$ to all the rest it suffices to check $\gamma_{\text{sy,op}}\lesssim\gamma_{\text{op}}$, since the reverse inequality can be verified to hold easily. Fix $\mu$ any admissible measure for $\gamma_{\text{{sy,op}}}(E)$ with $\gamma_{\text{{sy,op}}}(E)\leq 2\mu(E)$. Notice that since $\pazocal{P}_{\mu}$ and $\pazocal{P}^{\ast}_{\mu}$ are $n$-dimensional CZ operators, then $\pazocal{P}_{\text{sy},\mu}$ also satisfies such property. Therefore, following an analogous proof to [To3, Theorem 2.16] one deduces, from the $L^{2}(\mu)$-boundedness of $\pazocal{P}_{\text{sy},\mu}$, its boundedness as an operator from the space of Borel finite signed measures $M(\mathbb{R}^{n+1})$ to $L^{1,\infty}(\mu)$. That is, there exists a constant $C>0$ such that for any $\nu\in M(\mathbb{R}^{n+1})$,
@@ 3. Even more capacities
for all the pair of points where such expression makes sense. Notice that it is also a nonnegative kernel and differentiable, with respect to $\overline{x}$ or $\overline{y}$, out of a set of null Lebesgue measure. We may also define $P_{\Lambda}^{\ast}$ and $P_{\text{sy},\Lambda}$ in an analogous way. In fact, for the sake of notation, we shall prove three basic properties only for $P_{\Lambda}$, but they can be checked to hold also for $P_{\Lambda}^{\ast}$ and $P_{\text{sy},\Lambda}$ with almost exactly the same proofs. Essentially the convolution kernels $P_{\Lambda},P_{\Lambda}^{\ast}$ and $P_{\text{sy},\Lambda}$ in $\mathbb{R}^{n+1}$ define $n$-dimensional CZ operators. In fact, such suppressed kernels satisfy 3 as an additional property.
@@
Let $\Lambda:\mathbb{R}^{n+1}\to\mathbb{[}0,\infty)$ be a 1-Lipschitz function. Then, kernels $P_{\Lambda},P_{\Lambda}^{\ast}$ and $P_{\text{{sy}},\Lambda}$ define $n$-dimensional CZ operators which satisfy being “well suppressed” at the points where $\Lambda>0$. Namely, if we fix any $\overline{x}\neq\overline{y}$ points in $\mathbb{R}^{n+1}$, the precise estimates that $P_{\Lambda}$ satisfies ((as well as $P_{\Lambda}^{\ast}$ and $P_{\text{{sy}},\Lambda})$ are
@@
$\gamma_{\text{sy},2}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma_{n}(E),\,\|P_{\text{{sy}}}\ast\mu\|_{L^{2}(\mu)}\leq\mu(E)^{1/2}\big{\}}.$
It is clear that $\gamma_{\text{{sy}},\text{{op}}}(E)\leq\gamma_{\text{sy},2}(E)$. In the following lemma we prove that the reverse inequality also holds:
@@
Due to the above Lemma 3.2, we may add $\gamma_{\text{sy},2}$ to the statement of Theorem 2.1. In fact, the arguments above also prove that we can also add the following capacity to Theorem 2.1,
$\widetilde{\gamma}_{2}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma_{n}(E),\,\|P\ast\mu\|_{L^{2}(\mu)}\leq\mu(E)^{1/2},\|P^{\ast}\ast\mu\|_{L^{2}(\mu)}\leq\mu(E)^{1/2}\big{\}}.$
@@ 3.2. Normalization by a uniform bound at all points of the support
$\gamma^{\star}_{+}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma_{n}(E),\,|P\ast\mu(\overline{x})|\leq 1,\,\forall\overline{x}\in E\big{\}},$
@@
Indeed, given $\mu\in\Sigma_{n}(E)$ admissible for $\gamma_{\text{op}}(E)$ with $\gamma_{\text{op}}(E)\leq 2\mu(E)$, proceeding as in Remark 2.2 we can find a function $h:E\to[0,1]$ such that $\int_{E}h\text{d}\mu\geq\mu(E)/2$ and $\|\pazocal{P}(h\mu)\|_{\infty}\leq C$, for some $C>0$ constant. We know that the latter estimate implies $\|\pazocal{P}(h\mu)\|_{L^{\infty}(\mu)}\leq C^{\prime}$ for some other constant $C^{\prime}>0$ (by a Cotlar type inequality analogous to that of [MaP, Lemma 5.4]). Applying now Cotlar’s inequality of [To3, Theorem 2.18], for example, we have
@@ 3.3. An equivalent variational capacity
We shall consider a construction already presented in the proof of Lemma 3.2: let $\psi$ be a radial test function with $0\leq\psi\leq 1$, $\psi\equiv 0$ in $B_{1/2}(0)$, $\psi\equiv 1$ in $\mathbb{R}^{n+1}\setminus{B_{1}(0)}$ and such that $\|\nabla\psi\|_{\infty}\leq 1$. Fix $E\subset\mathbb{R}^{n+1}$ compact set and $\mu\in\Sigma_{n}(E)$. For each $\tau>0$, write $\psi_{\tau}(\cdot):=\psi(\cdot/\tau)$ and define, for each $\overline{x}\in\mathbb{R}^{n+1}$ and $f\in L^{1}_{\text{loc}}(\mu)$.
@@
Let us prove, using [MPr, Lemma 2.1] (also valid for $P^{\ast}$ and $P_{\text{{sy}}}$) , that the regularized (continuous) kernel $P_{\text{{sy}}}\psi_{\tau}$ defines a CZ convolution operator with constants not depending on $\tau$. It is clear that $|P_{\text{{sy}}}\psi_{\tau}(\overline{x})|\leq|\overline{x}|^{-n}$. Then, by the symmetry of $P_{\text{{sy}}}\psi_{\tau}$, it suffices to check that for any $\overline{x},\overline{x}^{\prime}\in\mathbb{R}^{n+1}$ with $\overline{x}\neq 0$ and $|\overline{x}-\overline{x}^{\prime}|\leq|\overline{x}|/2$,
@@
[…]
@@
[…]
@@
$\gamma^{\star}_{\text{sy},\psi_{\tau},+}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma_{n}(E),\,|P_{\text{{sy}}}\psi_{\tau}\ast\mu(\overline{x})|\leq 1,\,\forall\overline{x}\in E\big{\}}.$
@@
We begin by proving 1. Let $(\tau_{k})_{k}$ be a monotonically decreasing sequence to 0, and let $\mu_{k}\in\Sigma_{n}(E)$ be admissible for $\gamma^{\star}_{\text{{sy}},\psi_{\tau_{k}},+}(E)$ and such that $\gamma^{\star}_{\text{{sy}},\psi_{\tau_{k}},+}(E)\leq 2\mu_{k}(E)$. Passing to a subsequence if necessary, assume $\mu_{k}\rightharpoonup\mu_{0}$ weakly [Ma, Theorem 1.23], i.e.
@@
It is not hard to prove that $\mu_{0}\in\Sigma_{n}(E)$. Now fix $m>0$ integer and $\tau\in[\tau_{m+1},\tau_{m}]$, and also assume $k>m$. By (3.7) we get that for $\overline{x}\in E$,
@@
and 1 follows. To prove 2, take $\mu$ admissible for $\gamma^{\star}_{\text{sy},+}(E)$. Using (3.7),
@@
implying that $\gamma^{\star}_{\text{sy},\psi_{\tau},+}(E)\geq\frac{\mu(E)}{1+C}$, and by the arbitrariness of $\mu$ we finally deduce
$\inf_{\tau>0}\gamma^{\star}_{\text{sy},\psi_{\tau},+}(E)\gtrsim\gamma^{\star}_{\text{sy},+}(E).$
@@
$\gamma_{\text{sy},\psi_{\tau},+}^{\star}(E)\underset{(1)}{\leq}\gamma_{\text{sy},\psi_{\tau},2}(E)\underset{(2)}{\lesssim}\gamma_{\text{sy},\psi_{\tau},\text{op}}(E)\underset{(3)}{\lesssim}\gamma_{\text{sy},\psi_{\tau},+}(E)\underset{(4)}{\leq}\gamma_{\text{sy},\psi_{\tau},+}^{\star}(E).$
Inequality (1) is trivial. To verify (2) proceed as in the proof of $\gamma_{\text{sy},2}(E)\lesssim\gamma_{\text{sy},\text{op}}(E)$ in Lemma 3.2. The proof of (3) goes as in Remark 2.2, and it relies on [To3, Theorem 2.16] and [Ch, Ch.VII, Theorem 23]. Notice that the independence of the CZ constants of $P_{\text{{sy}}}\psi_{\tau}$ of $\tau$ is crucial to ensure the validity of (2) and (3). Finally, (4) holds by the continuity of $\pazocal{P}_{\text{sy},\psi_{\tau}}\mu$ for each $\tau>0$, that has already been argued in the proof of Lemma 3.2. Observe that the previous chain of estimates also holds changing every $\gamma_{\text{sy}}$ for $\widetilde{\gamma}$, where we ask the respective normalization conditions in each case for both kernels $P\psi_{\tau}$ and $P^{\ast}\psi_{\tau}$.
@@
Now, bearing in mind that we have found positive constants $C_{1},C_{2}$ so that
@@
we define the variational capacity as follows:
Definition 3.1.
Let $E\subset\mathbb{R}^{n+1}$ be a compact set and choose $\tau_{0}$ small enough so that $\gamma_{\text{{sy}},\psi_{\tau_{0}},2}(E)\leq 2C_{1}C_{2}\,\gamma_{\text{sy},\text{op}}(E).$ Let $\pazocal{S}$ be the convolution operator associated to the kernel $P_{\text{{sy}}}\psi_{\tau_{0}}$ (that depends on $\tau_{0}$ and thus on $E$). We define the variational capacity of $E$ as
$\gamma_{\text{var}}(E):=\sup\Bigg{\{}\frac{\mu(E)^{2}}{\mu(E)+\int_{E}|\pazocal{S}\mu|^{2}\mathop{}\!\mathrm{d}\mu}\,:\,\mu\in\Sigma_{n}(E)\Bigg{\}}.$
@@
For each $\mu\in\Sigma_{n}(E)$ define
@@
As it has already been mentioned in the proof of Lemma 3.3, it is not difficult to prove that $\Sigma_{n}(E)$ is closed and sequentially compact with respect to weak convergence of measures. In fact, $\Sigma_{n}(E)$ is compact (since it is contained in the space of finite signed Radon measures on $E$, which is metrizable, thought as the dual of the separable space $\pazocal{C}_{c}(\mathbb{R}^{n+1})$). Moreover, it is clear that if $\mu_{k}\rightharpoonup\mu_{0}$ on $\Sigma_{n}(E)$, then $\mu_{k}(E)\to\mu_{0}(E)$. In this setting, we also claim
@@
$\bigg{\rvert}$ $\int_{E}\bigg{(}\int_{E}f(\overline{x},\overline{y})\mathop{}\!\mathrm{d}\mu_{k}(\overline{x})\bigg{)}^{2}\mathop{}\!\mathrm{d}\mu_{k}(\overline{y})-\int_{E}\bigg{(}\int_{E}f(\overline{x},\overline{y})\mathop{}\!\mathrm{d}\mu_{0}(\overline{x})\bigg{)}^{2}\mathop{}\!\mathrm{d}\mu_{0}(\overline{y})\bigg{\rvert}$
@@
Further, observe that we can restrict ourselves to the case that there exists $\mu\in\Sigma_{n}(E)$ such that $F(\mu)>0$. If this was not the case, since $\int_{E}|\pazocal{S}\mu|\mathop{}\!\mathrm{d}\mu\lesssim\text{diam}(E)^{2n}/\tau_{0}^{n}<\infty$, we would have, necessarily, $\mu(E)=0$, implying, by definition, $\gamma_{\text{var}}(E)=0=\gamma_{\text{{sy}},\psi_{\tau_{0}},2}(E)$.
So let us assume that there exists $\mu\in\Sigma_{n}(E)$ such that $F(\mu)>0$, meaning $\gamma_{\text{var}}(E)>0$. Let us begin by proving that any extremal $\mu_{0}$ for $F$ in this setting satisfies
@@
If this was not the case, then $\int_{E}|\pazocal{S}\mu_{0}|^{2}\mathop{}\!\mathrm{d}\mu_{0}=M\mu_{0}(E)$ for some $M>1$. Define $\mu_{1}:=M^{-1/2}\mu_{0}\in\Sigma_{n}(E)$ and notice that
$F(\mu_{1})$ $=\frac{M^{-1}\mu_{0}(E)^{2}}{M^{-1/2}\mu_{0}(E)+\int_{E}|\pazocal{S}\mu_{1}|^{2}\mathop{}\!\mathrm{d}\mu_{1}}=\frac{\mu_{0}(E)}{2M^{1/2}}>\frac{\mu_{0}(E)}{1+M}=F(\mu_{0}),$
@@
Let us notice that (3.10) holds in general for any extremal measure $\mu_{0}$ for $F$. Indeed, in the case that for any $\mu\in\Sigma_{n}(E)$ one had $F(\mu)=0$, then $\mu(E)=0$. So any measure would be extremal and (3.10) would hold trivially.
Finally, we are ready to give an improved version of Theorem 2.1:
@@
For any $E\subset\mathbb{R}^{n+1}$ compact set,
$\gamma_{+}(E)\approx\gamma^{\star}_{+}(E),$
where the respective bounds of the potentials can be taken indifferently and independently for $P,P^{\ast},P_{\text{{sy}}}$ or both $P$ and $P^{\ast}$ simultaneously. Moreover, all the previous capacities are also comparable to $\gamma_{\text{{sy}},2}(E),\;\gamma_{\text{{op}}}(E),\;\gamma_{\text{{sy,op}}}(E)$ and $\gamma_{\text{{var}}}(E)$.